# The Rise and Fall of the Initial Era

### v1

@@ III.3 Geographical analysis
[…]
@@
Figure 9: Proportion of authors publishing their name using initial form rather than full form by country of affiliation from 1945 until present day. The grey dotted lines show all countries in the list for context, selected Western countries are picked out in colour.
@@
In Fig. 10 we have selected countries from the Asia-Pacific region and Brazil. In this case we see clearly the formalism (initial form alignment) of India from 1945 to present day with it being consistently (and in much of this time period significantly) above its comparators. Brazil passed the production threshold for this analysis in the late 1960s, at which point it was already more aligned to full name usage than most of the rest of the world (below most lines in the grey background), and continues to be even today. China, Japan and South Korea have all consistently been amongst the countries least likely to use initial form for the longest period of time. The facet in common for these countries is their use of non-Roman writing systems. Thus, in native language, author names are often single character, but when translated into English as choice is made to list the name as as a full name rather than just as a single initial. This may, in part, be due to cultural and linguistic factors in many Asian countries that give rise to common last names, the lack of a second or third given name, and hence the need to state full first names in order have the ability to disambiguate authors in a pre-ORCID environment Kretzschmar (2004); Velden et al. (2011); Liu et al. (2012). All these issues are compounded by transliteration in the scholarly record.
@@
Figure 12: Proportion of authors publishing with initial form in 2020 versus power distance for a selection of countries. The size of each disk represents the cumulative volume of publications associated with the country between 1945 and 2023. Each disk is coloured according to the linguistic grouping of the principal language spoken in the country of the author’s affiliation.
@@
In an attempt to gain further insight, we explore the interplay of linguistic and cultural effects further by turning to the work of the cultural theorist Hofstede from the late 1960s. Through his work at IBM, Hofstede developed Cultural Dimension theory Hofstede (1983, 2003, 2011) in an attempt to understand how cultural considerations affected organisational structures. One of the cultural dimensions that Hofstede developed, called the Power Distance Index, attempts to quantify the extent to which the less powerful members of organisations and institutions accept and expect power to be distributed unequally. As such, Power Distance may be thought of as a proxy for how formal a country is culturally in a professional setting. Of course, such things change with time and the Power Distance shown here is a snapshot at a particular point in time. We argue, however, that these things do not change as quickly from a cultural perspective as they do from a technological perspective.
@@ III.4 Disciplinary Analysis
Figures 14 and 15 show analogous plots to Figs 9, 10, and 11 but instead of breaking down the initial form percentage into country contributions, the basis has been changed to examine subject areas. In this case we make use of the 2020 ANZSRC Field of Research Codes that are automatically assigned to publications in Dimensions Porter et al. (2023). The plots must still, in aggregated form, reduce to Fig. 7 as was the case for the sum over all paths in the country-based plots and hence Fig. 13 shows the proportion of papers written in each of the subject areas and hence shows us the contribution level of each subject to the aggregate (or, the effective weighting of each line in each plot).
Figure 14: Proportion of authors using initial form on journal articles by year from 1945 to 2022. Selected Science, Technology, Engineering and Medicine (STEM) subjects are picked out in colour. Fields are determined via Dimensions automated attribution to ANZSRC 2020 coding.
@@
Figure 15: Proportion of authors using initial form on journal articles by year from 1945 to 2022. Selected Social Sciences, Humanities and the Arts for People and the Economy (SHAPE) subjects are picked out in colour. Fields are determined via Dimensions automated attribution to ANZSRC 2020 FoR coding.
@@
In C. P. Snow’s Tale of Two Cultures Snow (2012), he argues that Social Sciences, Arts and Humanities (which we will refer to as the SHAPE disciplines) have a different culture to Science, Technology, Engineering and Medicine (which we will refer to as the STEM disciplines), and this appears to be born out in our data.
@@ III.5 Journal analysis
In the mid nineties, Highwire press emerged as a significant player in the online hosting of Journal content. Starting with the Journal of Biological Chemistry, Highwire was at the forefront of re-imagining the definition of a paper from a digital ‘photocopy’ of a physical paper, Highwire’s technology allowed the paper to be a fully interactive digital object in a contemporary sense—text was textual rather than graphical and hence was searchable, elements of the body of the document were broken down into component parts so that images and their captions were distinct objects in the paper noa (2000); Watson (2011).
@@
The BMJ makes a similar shift, also in 1995, although its profile suggests that that first names are strongly preferred but not enforced Smith and Smith (1996). Prior to this, in around 1975, it appears that all three journals, updated their editorial policies as there are sharp changes of behaviour. Both the BMJ and the Journal of Biological Chemistry appear to have put journal policies in place that mandate (or close to mandate) initial form, while Tetrahedron appears to have either removed a mandate for initial form or even actively promoted full name form given the initial rapid drop off from close to 100% to 60% initial-form name usage within just five years from 1975-1980.
@@ III.6 Technological Analysis
Figure 17 shows distinct and different behaviours between authors associated with articles appearing on PubMed that are associated with DOIs and those which aren’t. The blue line shows the percentage of authors associated with a DOI where initial form is used; the red line shows the same percentage for authors associated with articles without a DOI. For articles with a DOI, we see a gradual change to lower rates of initial form being used with authors migrating to full form. We note a small drop in the blue line around 2002 which, we speculate, would be associated with the launch of PubMed’s new platform. This new platform supported given name metadata fields and hence naturally encouraged the use of full form names NLM (2002, 2009).
Articles without a DOI (only having a PubMed identifier), shown in the red line of Fig. 17, have a distinctly different shape. The difference in behaviour of the red line is explained by a systemic development: When Crossref introduced the opportunity to implement DOIs for academic articles, their metadata format was sufficiently fully formed and modern to allow for a field to include given name and publishers adopting the new standard were able to update their back catalogues with the additional data. Thus, the blue line is a representation of the actual metadata included on articles as they were published, whereas the red line is the result of the legacy metadata standard that existed prior to the advent of Crossref’s metadata schema Crossref (2009). Even though articles without a DOI may, in fact, have full form names on the publication this is not represented in the metadata.
Similar parallels in behaviour can be seen with other major bibliographic and bibliometric databases that existed through this period, with Web of Science making collecting first names from records processed after 2007, and allowing them to be searched in 2011 Clarivate (2022), which Scopus had done in 2004 Hane (2004); Sullo (2007).
The ongoing decline in the use of initial form from 2002 to present day, may not be due only to technology changes but rather may be motivated by changes to another scholarly institution—research evaluation. In the period from 2002, technological developments took place to introduce of (current research information systems) CRISes Jörg (2010), to address the need for reporting data in research economies that introduced research evaluation such as the UK and Australia. It was not until 2010 the ORCID launched formally to tackle the disambiguation problem at a deeper systemic level - perhaps surprisingly the launch of ORCID does not seem to have a particularly direct effect on these data Butler (2012); Schiermeier (2015); Meadows et al. (2019).
@@ III.7 Gender Analysis
Finally, we turn our attention towards a gender-focused analyses. We have pointed out reasons that gender may be hidden by choice in certain contexts in the introduction to this paper. However, we believe that the data that we can show as a result of analyses in this paper demonstrates, perhaps unsurprisingly, that technology can be a significant modifier to gender visibility in publication. We note that the results shown in this section rely on statistical methods and hence have known limitations, however, we do believe them to be sufficiently robust for the purposes of a high-level indicative analysis, as presented here. The work of Lockhart et al. provides a critical understanding of some of the issues with this type of analysis Lockhart et al. (2023).
@@
While the reporting of gender participation in research is still relatively nascent with systematised aggregation of reporting only available since around 1996. Already in many regions of the world that reported in 2002, the date of the change in PubMed systems, women regularly accounted for between 30% and 50% of researchers for Statistics . Yet, PubMed authorships at the time attribute only 15% of output to those with names statistically associated with women, while around 32% of output was either of indeterminate or unknown gender. This, together with the steeper rise in participation of women in using full form names, suggests that the initial form convention proportionally over represented women over men, making the contributions of women researchers systematically more difficult to identify.
@@ IV.1 The rise of the Initial Era
We speculate that the adoption of initial form during this exceptional 40 year period is far from Vannevar Bush’s Endless Frontier Bush (1945) being accidental, as it was a highly influential document that appeared at the dawn of US-preeminence, a time when science was being established as “serious business” with a more formal cultural norm associated with it. This may have been influenced by the rise in importance of science and medicine and the view that technology could solve all problems. Similarly, the Marshall Plan rebuilding of Europe could have led to the import of norms the US, but this seems less likely since the US is continually one of the lower users of initial-form (see Fig. 9).
However, during this period, research output was dominated by the US, the UK and Japan—all countries that were predisposed to publish in English and to reflect the cultural norms in the US for political reasons. The Bretton-Woods Era came to an end in the late 1970s and the Cold War in the late 1980s, and with the rise of globalisation greater international collaboration has been possible, leading to Adam’s Fourth Age Adams (2013). Yet, as we see from Figs. 9, 10 and 11, the use of initial form persisted (and even increased, possibly due to the constraints of paper media and their interaction with the era of Big Science) from 1980 to 2000. As demonstrated in our technological analysis, it was a combination of house style decisions, technology changes in the form of PubMed and Crossref, and the switch to online journal submission systems that fundamentally changed research culture back to full name usage. It remains a “chicken and egg” issue as to whether technology evolved to meet cultural needs or whether technological change gave cultural norms an opportunity to assert themselves.
As collaboration has become more of a norm in research (see Fig. 2) and as many countries have invested in and developed their own advanced research economies (see Fig. 13), a much more diverse community has emerged. This will necessarily require the representation of newly evolved norms. Yet, the data landscape to explore the emergence of these norms is a challenging one.
@@ IV.2 The fall of the Initial Era
The articulation of this need for openness is often found in evaluation mechanisms, either national or funder based. All the factors that we identify above relate to evaluation. Technological developments, especially around name disambiguation, relate to the rise of ORCID as a standard Haak et al. (2018); Porter (2022), the use research information management systems and research profiling tools Ilik et al. (2018). The globalisation of research has led to increased researcher movement and for researchers to be able to move freely, they need to be able to assert their credentials—which involves asserting identity in relation to their research work. And, in a world where nefarious actors are on the rise Porter and McIntosh (2024), understanding the provenance of research often relies on understanding identity.
@@ IV.3 The future of bibliometric archaeology
If we have learned anything in the writing of this paper, we must reflect that the lens through which we look is as important as the analysis itself. While Dimensions is a useful tool in understanding demographics and evolution of certain types of literature, it is a product of its world and of its construction Hook et al. (2018). Data coverage pre-1800 shows low volumes of output and, in part, this is because output levels are lower but it is also because strategic choices have had to be made by infrastructure organisations in what should be afforded, the cost and effort of creating digital versions and digital identifiers for past material that is not likely to be of current research value in a non-sociohistorical setting. This mirrors the more hidden lack of data that we have identified in the Initial Era, where the contribution of women is hidden through the technical implementation of a social construction.
@@
It is apposite to ask whether we live in a special time in bibliometric analysis where name data continues to be freely available and whether this period will persist. ORCID now assigns unique identifiers to any researcher who would like to have one. We live in a time of increasing polarisation and researchers who wish to carry out certain types of research or express particular opinions can face censure or removal of funding. If we take the logic of research identifiers further, some might argue that the adoption of cryptographic identities, like Satoshi Nakamoto, would allow for researchers to regain intellectual freedom. Zero knowledge proofs could be used to authenticate claims for career progression, double-blind peer review would be built in to such a system, potentially increasing the overall fairness of both publication and funding practices. Others might argue that this undermines deeply the trust that we have in the scholarly record and permits abusive expression or propagation of fake research or support of political or commercial interests from “behind a mask”.
@@
Within contribution statements that use the Credit Ontology Allen et al. (2014), initials are also commonly employed to refer to authors although this is not part of the standard. This convention also creates disambiguation issues when two authors share the same surname and first initials. Here too, as the Digital Structure of a paper continues to evolve, we should be careful not to unquestioningly embed the naming conventions of a different era into our evolving metadata standards.

### v2

@@ III.3 Geographical analysis
The data in Fig. 7 are averaged across all countries but, over the time period that we are examining, not all countries participate in the research enterprise equally. Thus there is an implicit weighting in the global picture toward specific countries. If we speak about cultural, political, linguistic or geographic trends then it is important to understand the level of participation of different countries to an international average. As such, Fig. 8 reveals the contribution of each country to global paper production rates in Dimensions. The graph is produced by taking each paper and partitioning it by the countries of the affiliations of each author on the paper and then summing these contributions over all journal articles published in each year. In the plot, we have only plotted the top 13 countries and have conglomerated under the title ”Rest of World” as a 14th participant. As we can see from the Figure, during the period from 1945 to 1990, an extremely high percentage of global research output emanated from the US and the UK. While Japan and Germany make a significant and sustained contribution from the mid-1950s, between 50% and 60% of global output in the 1970s and 1980s came from the US and the UK - English-speaking countries with a somewhat shared cultural base. Thus, the behaviours summarised in Fig. 7 are dominated by an Anglo-American societal behaviours. However, as we see in Fig. 9 cultural alignment between the UK and the US may be more divergent than expected.
@@
Figure 9: Proportion of papers only displaying initial-form names broken down by countries of author affiliations from 1945 until present day. The grey dotted lines show all countries in the list for context, selected Western countries are picked out in colour. Fractional apportionment of papers to countries has been applied.
@@
In Fig. 10 we have selected countries from the Asia-Pacific region and Brazil. In this case we see clearly the formalism (initial form alignment) of India from 1945 to present day with it being consistently (and in much of this time period significantly) above its comparators. Brazil passed the production threshold for this analysis in the late 1960s, at which point it was already more aligned to full name usage than most of the rest of the world (below most lines in the grey background), and continues to be even today. China, Japan and South Korea have all consistently been amongst the countries least likely to use initial form for the longest period of time. The facet in common for these countries is their use of non-Roman writing systems. Thus, in native language, author names are often single character, but when translated into English as choice is made to list the name as as a full name rather than just as a single initial. This may, in part, be due to cultural and linguistic factors in many Asian countries that give rise to common last names, the lack of a second or third given name, and hence the need to state full first names in order have the ability to disambiguate authors in a pre-ORCID environment [31, 32, 33]. All these issues are compounded by transliteration in the scholarly record.
@@
Figure 12: Proportion of papers published displaying only initial-form authors in 2020 versus power distance for a selection of countries. The size of each disk represents the cumulative volume of publications associated with the country between 1945 and 2023. Each disk is coloured according to the linguistic grouping of the principal language spoken in the country of the author’s affiliation.
@@
In an attempt to gain further insight, we explore the interplay of linguistic and cultural effects further by turning to the work of the cultural theorist Hofstede from the late 1960s. Through his work at IBM, Hofstede developed Cultural Dimension theory [34, 35, 36] in an attempt to understand how cultural considerations affected organisational structures. One of the cultural dimensions that Hofstede developed, called the Power Distance Index, attempts to quantify the extent to which the less powerful members of organisations and institutions accept and expect power to be distributed unequally. As such, Power Distance may be thought of as a proxy for how formal a country is culturally in a professional setting. Of course, such things change with time and the Power Distance shown here is a snapshot at a particular point in time. We argue, however, that these things do not change as quickly from a cultural perspective as they do from a technological perspective.
@@ III.4 Disciplinary Analysis
Figures 14 and 15 show analogous plots to Figs 9, 10, and 11 but instead of breaking down the initial form percentage into country contributions, the basis has been changed to examine subject areas. In this case we make use of the 2020 ANZSRC Field of Research Codes that are automatically assigned to publications in Dimensions [37]. The plots must still, in aggregated form, reduce to Fig. 7 as was the case for the sum over all paths in the country-based plots and hence Fig. 13 shows the proportion of papers written in each of the subject areas and hence shows us the contribution level of each subject to the aggregate (or, the effective weighting of each line in each plot).
Figure 14: Proportion of journal articles in which authors only use initial form assigned to ANZSRC FoR-subject by year from 1945 to 2022. Selected Science, Technology, Engineering and Medicine (STEM) subjects are picked out in colour. Fractional apportionment is not applied - there is duplication of counting if papers cross 2-digit-FoR classifications. Fields are determined via Dimensions automated attribution to ANZSRC 2020 coding.
@@
Figure 15: Proportion of journal articles in which authors only use initial form assigned to ANZSRC FoR-subject by year from 1945 to 2022. Selected Social Sciences, Humanities and the Arts for People and the Economy (SHAPE) subjects are picked out in colour. Fractional apportionment is not applied - there is duplication of counting if papers cross 2-digit-FoR classifications. Fields are determined via Dimensions automated attribution to ANZSRC 2020 FoR coding.
@@
In C. P. Snow’s Tale of Two Cultures [38], he argues that Social Sciences, Arts and Humanities (which we will refer to as the SHAPE disciplines) have a different culture to Science, Technology, Engineering and Medicine (which we will refer to as the STEM disciplines), and this appears to be born out in our data.
@@ III.5 Journal analysis
In the mid nineties, Highwire press emerged as a significant player in the online hosting of Journal content. Starting with the Journal of Biological Chemistry, Highwire was at the forefront of re-imagining the definition of a paper from a digital ‘photocopy’ of a physical paper, Highwire’s technology allowed the paper to be a fully interactive digital object in a contemporary sense—text was textual rather than graphical and hence was searchable, elements of the body of the document were broken down into component parts so that images and their captions were distinct objects in the paper [39, 40].
@@
The BMJ makes a similar shift, also in 1995, although its profile suggests that that first names are strongly preferred but not enforced [41]. Prior to this, in around 1975, it appears that all three journals, updated their editorial policies as there are sharp changes of behaviour. Both the BMJ and the Journal of Biological Chemistry appear to have put journal policies in place that mandate (or close to mandate) initial form, while Tetrahedron appears to have either removed a mandate for initial form or even actively promoted full name form given the initial rapid drop off from close to 100% to 60% initial-form name usage within just five years from 1975-1980.
@@ III.6 Technological Analysis
Figure 17 shows distinct and different behaviours between authors associated with articles appearing on PubMed that are associated with DOIs and those which aren’t. The blue line shows the percentage of authors associated with a DOI where initial form is used; the red line shows the same percentage for authors associated with articles without a DOI. For articles with a DOI, we see a gradual change to lower rates of initial form being used with authors migrating to full form. We note a small drop in the blue line around 2002 which, we speculate, would be associated with the launch of PubMed’s new platform. This new platform supported given name metadata fields and hence naturally encouraged the use of full form names [42, 43].
Articles without a DOI (only having a PubMed identifier), shown in the red line of Fig. 17, have a distinctly different shape. The difference in behaviour of the red line is explained by a systemic development: When Crossref introduced the opportunity to implement DOIs for academic articles, their metadata format was sufficiently fully formed and modern to allow for a field to include given name and publishers adopting the new standard were able to update their back catalogues with the additional data. Thus, the blue line is a representation of the actual metadata included on articles as they were published, whereas the red line is the result of the legacy metadata standard that existed prior to the advent of Crossref’s metadata schema [44]. Even though articles without a DOI may, in fact, have full form names on the publication this is not represented in the metadata.
Similar parallels in behaviour can be seen with other major bibliographic and bibliometric databases that existed through this period, with Web of Science making collecting first names from records processed after 2007, and allowing them to be searched in 2011 [45], which Scopus had done in 2004 [46, 47].
The ongoing decline in the use of initial form from 2002 to present day, may not be due only to technology changes but rather may be motivated by changes to another scholarly institution—research evaluation. In the period from 2002, technological developments took place to introduce of (current research information systems) CRISes [48], to address the need for reporting data in research economies that introduced research evaluation such as the UK and Australia. It was not until 2010 the ORCID launched formally to tackle the disambiguation problem at a deeper systemic level - perhaps surprisingly the launch of ORCID does not seem to have a particularly direct effect on these data [49, 50, 51].
@@ III.7 Gender Analysis
Finally, we turn our attention towards a gender-focused analyses. We have pointed out reasons that gender may be hidden by choice in certain contexts in the introduction to this paper. However, it is clear that the reasons for the obscuring of gender in research publication are considerably more subtle [52, 53]. Countless studies demonstrate differences in outcome for women in peer review in both grant and publication contexts, and citations of their work is lower than for male counterparts [54, 55, 56, 28]. Further work seeks to understand the causal nature of these effects [57]. What we show here is that our data analysis as a further dimension for consideration namely that, perhaps unsurprisingly, technology can be a significant modifier to gender visibility in publication. We note that the results shown in this section rely on statistical methods and hence have known limitations, however, we do believe them to be sufficiently robust for the purposes of a high-level indicative analysis, as presented here. The work of Lockhart et al.[58] provides a critical understanding of some of the technical issues with this type of analysis.
@@
While the reporting of gender participation in research is still relatively nascent with systematised aggregation of reporting only available since around 1996. Already in many regions of the world that reported in 2002, the date of the change in PubMed systems, women regularly accounted for between 30% and 50% of researchers [59]. Yet, PubMed authorships at the time attribute only 15% of output to those with names statistically associated with women, while around 32% of output was either of indeterminate or unknown gender. This, together with the steeper rise in participation of women in using full form names, suggests that the initial form convention proportionally over represented women over men, making the contributions of women researchers systematically more difficult to identify.
@@ IV.1 The rise of the Initial Era
We speculate that the adoption of initial form during this exceptional 40-year period is far from Vannevar Bush’s Endless Frontier [24] being accidental, as it was a highly influential document that appeared at the dawn of US-preeminence, a time when science was being established as “serious business” with a more formal cultural norm associated with it. This may have been influenced by the rise in importance of science and medicine and the view that technology could solve all problems. Similarly, the Marshall Plan rebuilding of Europe could have led to the import of norms the US, but this seems less likely since the US is continually one of the lower users of initial-form (see Fig. 9).
However, during this period, research output was dominated by the US, the UK and Japan—all countries that were predisposed to publish in English and to reflect the cultural norms in the US for political reasons. The Bretton-Woods Era came to an end in the late 1970s and the Cold War in the late 1980s, and with the rise of globalisation greater international collaboration has been possible, leading to Adam’s Fourth Age [19]. Yet, as we see from Figs. 9, 10 and 11, the use of initial form persisted (and even increased, possibly due to the constraints of paper media and their interaction with the era of Big Science) from 1980 to 2000. As demonstrated in our technological analysis, it was a combination of house style decisions, technology changes in the form of PubMed and Crossref, and the switch to online journal submission systems that fundamentally changed research culture back to full name usage. It remains a “chicken and egg” issue as to whether technology evolved to meet cultural needs or whether technological change gave cultural norms an opportunity to assert themselves.
The methodology behind our analysis in Sec. III.3 is prone to reflecting international trends. By construction, the unit of analysis is the paper and is classified as an initial-form paper only in the situation that all authors use that form. There are normative trends implicit in a global research community—publishers have house styles that are applied regardless of origin; if the authors who brought funding for the work, or who are in some other way senior, choose a style of publishing their name on the paper perhaps there is a psychological bias that goes on that we cannot track in the data; if a majority of authors choose a specific style then how many authors are comfortable breaking step and choosing a different form for their name? International collaboration is another anchor to which our methodology is particularly sensitive as only papers that are entirely initial-based factor in analysis such as Sec. III.3—hence, a core assumption of the analysis is that the kinds of effects discussed here were sufficiently strong as to normalise behaviour across a paper regardless of geographic background.
As collaboration has become more of a norm in research (see Fig. 3) and as many countries have invested in and developed their own advanced research economies (see Fig. 13), a much more diverse community has emerged. Hence, once “senior partners” are now just simple partners. On the one hand, this will necessarily require the representation of newly evolved norms. Yet, the norms are already set and new norms take time to evolve.
@@ IV.2 The fall of the Initial Era
The articulation of this need for openness is often found in evaluation mechanisms, either national or funder based. All the factors that we identify above relate to evaluation. Technological developments, especially around name disambiguation, relate to the rise of ORCID as a standard [60, 61], the use research information management systems and research profiling tools [62]. The globalisation of research has led to increased researcher movement and for researchers to be able to move freely, they need to be able to assert their credentials—which involves asserting identity in relation to their research work. And, in a world where nefarious actors are on the rise [63], understanding the provenance of research often relies on understanding identity.
@@ IV.3 The future of bibliometric archaeology
If we have learned anything in the writing of this paper, we must reflect that the lens through which we look is as important as the analysis itself. While Dimensions is a useful tool in understanding demographics and evolution of certain types of literature, it is a product of its world and of its construction [30]–specifically the requirement of a DOI or other mainstream identifier limits Dimensions’ coverage at this time. Data coverage pre-1800 shows low volumes of output and, in part, this is because output levels are lower but it is also because strategic choices have had to be made by infrastructure organisations in what should be afforded, the cost and effort of creating digital versions and digital identifiers for past material that is not likely to be of current research value in a non-sociohistorical setting. This mirrors the more hidden lack of data that we have identified in the Initial Era, where the contribution of women is hidden through the technical implementation of a social construction.
@@
It is apposite to ask whether we live in a special time in bibliometric analysis where name data continues to be freely available and whether this period will persist. ORCID now assigns unique identifiers to any researcher who would like to have one. We live in a time of increasing polarisation and researchers who wish to carry out certain types of research or express particular opinions can face censure or removal of funding. If we take the logic of using research identifiers further, some might argue that the adoption of cryptographic identities, like Satoshi Nakamoto, would allow for researchers to regain intellectual freedom. Zero-knowledge proofs could be used to authenticate claims for career progression, double-blind peer review would be built in to such a system, potentially increasing the overall fairness of both publication and funding practices. Others might argue that this undermines deeply the trust that we have in the scholarly record and permits abusive expression or propagation of fake research or support of political or commercial interests from “behind a mask”. This is, perhaps, the modern-day equivalent of the binaries pointed out by Evans et al. [64] in their commentary on the role of the algorithm.
@@
Within contribution statements that use the Credit Ontology [65], initials are also commonly employed to refer to authors although this is not part of the standard. This convention also creates disambiguation issues when two authors share the same surname and first initials. Here too, as the Digital Structure of a paper continues to evolve, we should be careful not to unquestioningly embed the naming conventions of a different era into our evolving metadata standards.
V Acknowledgements
The authors wish to thank Briony Fane for her careful reading and commenting on the manuscript, and to Kathryn Weber Boer for her suggestions to improve discussions on issues of gender.
VI Data Availability
The data and code for this paper is available from Figshare at: https://doi.org/10.6084/m9.figshare.25664154
VII Author Contribution
Simon J Porter: Conceptualization; Formal Analysis; Methodology; Visualisation; Writing - original draft; Writing - review & editing. Daniel W Hook: Formal Analysis; Methodology; Visualisation; Writing - original draft; Writing - review & editing.
VIII Conflict of interests
The authors of this paper are both employees of Digital Science, the owner and operator of Dimensions.